# Superconvergent and Divergence-Free Finite Element Methods for Stokes Equation

### v1

@@ Abstract.
Superconvergent and divergence-free finite element methods for the Stokes equation are developed. The velocity and pressure are discretized using $H(\mathrm{div})$-conforming vector elements and discontinuous piecewise polynomials. The discrete formulation employs a weak deviatoric gradient operator built with tangential-normal continuous finite elements for traceless tensors, requiring no stabilization. Optimal and superconvergent error estimates are established. The method connects to nonconforming virtual element and pseudostress-velocity-pressure mixed formulations. Numerical experiments verify the theory.
@@ 1. Introduction
In this work, we develop superconvergent and divergence-free finite element methods for the Stokes equation on a bounded domain $\Omega\subset\mathbb{R}^{d}$ ($d\geq 2$) with external force $\boldsymbol{f}\in L^{2}(\Omega;\mathbb{R}^{d})$:
(1) $\begin{cases}-\Delta\boldsymbol{u}-\nabla p=\boldsymbol{f}&\text{in }\Omega,\\ \operatorname{div}\boldsymbol{u}=0&\text{in }\Omega,\\ \boldsymbol{u}=0&\text{on }\partial\Omega,\end{cases}$
@@
where $\boldsymbol{u}$ and $p$ denote the velocity and pressure.
@@
Classical stable finite element pairs for (1) include the Taylor–Hood element [50], the MINI element [3], and the nonconforming $P_{1}$$P_{0}$ element [25]. More examples and a detailed discussion are given in [10]. However, these classical pairs are not pointwise divergence-free, which weakens mass conservation and reduces pressure robustness in error estimates. A review of the importance of divergence-free constraints and their numerical implications can be found in [41].
@@
Divergence-free finite element pairs include the Scott–Vogelius element [48], which requires special mesh conditions; smooth finite element pairs [56, 17, 18] with supersmooth degrees of freedom; conforming pairs on split meshes [6, 55, 22, 28, 32, 38]; conforming rational pairs [33, 34]; and low-order, normal-continuous nonconforming pairs [43, 49, 54]. We also refer to [8, 52, 51] for discontinuous Galerkin methods, and to [9, 20, 53, 40] for virtual element methods.
One can impose the divergence-free condition by using an $H(\operatorname{div})$-conforming finite element space for the velocity. Specifically, we use the Raviart–Thomas (RT) element space [47, 45] or the Brezzi–Douglas–Marini (BDM) element space [12, 46]:
@@
$\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}:=\{\boldsymbol{v}_{h}\in H_{0}(\operatorname{div},\Omega):\boldsymbol{v}_{h}|_{T}\in\mathbb{P}_{k}(T;\mathbb{R}^{d})+\mathbb{H}_{\ell}(T)\boldsymbol{x}\;\text{ for }T\in\mathcal{T}_{h}\}$
for the velocity, with integers $\ell=k$ (${\rm RT}_{k}$) or $\ell=k-1$ (${\rm BDM}_{k}$), and the discontinuous piecewise polynomial space $\mathbb{P}_{\ell}(\mathcal{T}_{h})$ for the pressure.
Since $\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\not\subset H^{1}(\Omega;\mathbb{R}^{d})$, the discretization of $\Delta\boldsymbol{u}$ is nontrivial. One approach is to use discontinuous Galerkin (DG) methods for the Laplacian operator, as in [52, 23]. The DG method in [52] introduces a penalty term to ensure stability, while the hybridizable DG (HDG) method in [23] employs $k$th-order polynomial spaces for all variables. In fact, it suffices to take $p_{h}\in\mathbb{P}_{k-1}(\mathcal{T}_{h})$ when the velocity $\boldsymbol{u}_{h}\in{\rm BDM}_{k}$.
In the vorticity–velocity–pressure formulation [26, 27], the vorticity $\omega={\rm curl\,}\boldsymbol{u}$ is introduced, and
$\Delta\boldsymbol{u}={\rm curl\,}\omega-{\rm grad\,}\operatorname{div}\boldsymbol{u}={\rm curl\,}{\rm curl\,}\boldsymbol{u}-{\rm grad\,}\operatorname{div}\boldsymbol{u}$
is interpreted as the Hodge Laplacian. However, as observed in [26, 27], for no-slip Dirichlet boundary conditions ($\boldsymbol{u}=0$ on $\partial\Omega$), there are stability issues and reduced convergence rates. This was later explained in [4]: the suboptimal convergence rate comes from the lack of a proper Hilbert complex in this setting.
The lost convergence rate can be partly recovered by mesh symmetry. In the triangular MAC (TMAC) schemes for the Stokes equation [21], mass lumping is applied to the vorticity so that it can be eliminated, and the velocity and pressure convergence rates are improved by mesh symmetry. Using the element pair $\textrm{RT}_{0}$$\textrm{P}_{0}$ for velocity and pressure (i.e., $k=\ell=0$) yields
$|I_{0,0}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}|_{1,h}+\|Q_{0}p-p_{h}\|\lesssim h^{\min\{1,s\}}|\ln h|^{1/2},$
where $s\geq\tfrac{1}{2}$ measures mesh symmetry and $I^{\rm\operatorname{div}}_{k,\ell}$ is the interpolation operator into $\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$. With the pair $\textrm{BDM}_{1}$$\textrm{P}_{0}$ (i.e., $k=1$, $\ell=0$), the TMAC scheme achieves
$|I_{1,0}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}|_{1,h}+\|Q_{0}p-p_{h}\|\lesssim h^{1+\min\{1/2,s\}}.$
In [30, 31], the mass-conserving mixed formulation with stresses (MCS) finite element method was developed for the pseudostress–velocity–pressure formulation, also employing $H(\operatorname{div})$-conforming velocity spaces. The stress space in [31] is slightly nonconforming, leading to the standard $h^{k}$ convergence instead of the superconvergent $h^{k+1}$ rate. In [30], a weak symmetry constraint was introduced and the stress space was enriched, yielding $h^{k+1}$ convergence for the $\mathrm{RT}_{k}$$\mathrm{P}_{k}$ pair with $k\geq 1$.
A divergence free weak virtual element method was developed in [20], which requires stabilization and does not yield superconvergence. The main idea is to introduce a piecewise vector-valued polynomial $\boldsymbol{\lambda}_{h}\in\mathring{\Lambda}_{k}:=\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h};\mathbb{R}^{d-1})$ defined on faces to enforce tangential continuity. We follow this approach and improve it to a stabilization-free scheme with superconvergence. Define the weak deviatoric gradient operator element-wise by
$(\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{v},\boldsymbol{\mu}),\boldsymbol{\tau})_{T}=-(\boldsymbol{v},\operatorname{div}\boldsymbol{\tau})_{T}+(\boldsymbol{n}\cdot\boldsymbol{v},\boldsymbol{n}^{\intercal}\boldsymbol{\tau}\boldsymbol{n})_{\partial T}+(\boldsymbol{\mu},\Pi_{F}\boldsymbol{\tau}\boldsymbol{n})_{\partial T}$
for any $\boldsymbol{\tau}\in\mathbb{P}_{k}(T;\mathbb{T})$ and $T\in\mathcal{T}_{h}$. The tensor space is reduced to the traceless tensor $\mathbb{T}$, since for $\boldsymbol{\sigma}:={\rm grad\,}\boldsymbol{u}$ with $\operatorname{div}\boldsymbol{u}=0$, we have $\operatorname{tr}\boldsymbol{\sigma}=\operatorname{tr}{\rm grad\,}\boldsymbol{u}=\operatorname{div}\boldsymbol{u}=0$.
We propose the following mixed finite element method: find $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$, $\boldsymbol{\lambda}_{h}\in\mathring{\Lambda}_{k}$, and $p_{h}\in\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R}$ with integers $k\geq 0$ and $\ell\in\{k,k-1\}$, such that
(2b) $(\operatorname{div}\boldsymbol{u}_{h},q_{h})$ $=0$
for all $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$, $\boldsymbol{\mu}_{h}\in\mathring{\Lambda}_{k}$, and $q_{h}\in\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R}$. Notably, the discrete method (2) requires no additional stabilization, despite relying on the weak differential operator $\operatorname{dev}{\rm grad\,}_{w}$.
We derive superconvergent estimates for the velocity and pressure:
$\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|+\|\operatorname{dev}{\rm grad\,}_{w}(I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h})\|+\|Q_{\ell}p-p_{h}\|\lesssim h^{k+1}|\boldsymbol{\sigma}|_{k+1},$
where $\boldsymbol{\sigma}=\operatorname{dev}{\rm grad\,}\boldsymbol{u},$ and $\boldsymbol{\sigma}_{h}:=\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h})$. Unlike the TMAC scheme [21], such superconvergence does not require the symmetry of the triangulation. It is further utilized to construct a new superconvergent approximation of $\boldsymbol{u}$.
Our method allows $\ell=0$ an important case not covered by MCS [30, 31]. We provide an illustration of the case $(k,\ell)=(0,0)$, which can be treated as a generalization of the classical MAC scheme [35] on rectangular grids to triangulations and achieve first order $h$ convergence, no loss of convergence compared with the vorticity-velocity-pressure formulation. One can use $\textrm{BDM}_{1}$$\textrm{P}_{0}$ pair, $(k,\ell)=(1,0)$, to get a second order $h^{2}$ scheme.
@@
We propose a hybridization of (2) by relaxing the normal continuity of the velocity, which enables the choice $(k,\ell)=(0,-1)$. The velocity is approximated by the piecewise constant vector space, and the normal continuity is enforced through a nonconforming linear element space for the pressure $p_{h}$. Even in this low-order setting, the method achieves at least first-order convergence, which is a notable feature.
The remainder of this paper is organized as follows. Section 2 discusses the weak div stability and the distributional discretization of the vector Laplacian. Section 3 develops mixed finite element methods for the Stokes equation, and Section 4 presents their equivalent discrete formulations. Finally, Section 5 reports numerical experiments that confirm and illustrate the theoretical results.

### v2

# Superconvergent and Divergence-Free Mixed Finite Element Methods for The Stokes Equation

@@ Abstract.
This paper develops divergence-free mixed finite element methods for the Stokes equation. Using H(div)-conforming velocities and discontinuous pressures ensures the inf-sup condition for the velocity–pressure pair and yields pointwise divergence-free velocities. However, this choice makes the vector Laplacian difficult to discretize. Inspired by mass-conserving mixed formulations with stresses, tangential–normal continuous traceless tensor elements are introduced to discretize the vector Laplacian. An inf-sup condition for the weak div operator between the stress and velocity spaces is then proved. Two key properties characterize the scheme. First, the stress–velocity inf-sup stability gives a stable discretization of the vector Laplacian without additional stabilization, unlike discontinuous Galerkin or virtual element methods. Second, the scheme has the property that if a stress field is weakly divergence-free, then it is also strongly divergence-free. This decouples the stress and velocity errors and leads to superconvergence. As a result, optimal-order error estimates are obtained for the stress, while the velocity and pressure converge at rates higher than the approximation orders of the chosen spaces. Numerical experiments confirm the theoretical results.
@@ 1. Introduction
In this work, we develop mixed finite element methods for the Stokes equations on a bounded domain $\Omega\subset\mathbb{R}^{d}$ ($d\geq 2$) with a given body force $\boldsymbol{f}\in L^{2}(\Omega;\mathbb{R}^{d})$:
(1) $-\Delta\boldsymbol{u}+\nabla p$ $=\boldsymbol{f}$ $\text{in }\Omega,$
$\operatorname{div}\boldsymbol{u}$ $=0$ $\text{in }\Omega,$
$\boldsymbol{u}$ $=\boldsymbol{0}$ $\text{on }\partial\Omega,$
@@
where $\boldsymbol{u}\in H_{0}^{1}(\Omega;\mathbb{R}^{d})$ and $p\in L^{2}_{0}(\Omega)$ represent the velocity and pressure, respectively.
Standard velocity–pressure finite element schemes for the Stokes equations depend crucially on the discrete divergence operator induced by the bilinear form $(\operatorname{div}\boldsymbol{v}_{h},q_{h})$. A stable scheme requires the discrete spaces $\boldsymbol{V}_{h}$ and $P_{h}$ to satisfy the inf-sup (or LBB) condition [12]:
(2) $\inf_{q_{h}\in P_{h}\setminus\{0\}}\sup_{\boldsymbol{v}_{h}\in\boldsymbol{V}_{h}\setminus\{\boldsymbol{0}\}}\frac{(\operatorname{div}\boldsymbol{v}_{h},q_{h})}{\|\boldsymbol{v}_{h}\|_{\boldsymbol{V}_{h}}\|q_{h}\|_{P_{h}}}\geq\beta>0,$
@@
where $\beta$ is a constant independent of the mesh size. Standard stable pairs for (1) include the Taylor–Hood element [52], the MINI element [3], and the nonconforming $P_{1}$$P_{0}$ element [26]; a comprehensive review is available in [10]. However, these elements do not enforce the divergence-free condition pointwise. Losing exact mass conservation often leads to poor pressure robustness in the velocity error estimates. We refer to [42] for a detailed discussion on the importance of strong divergence-free constraints.
Enforcing the discrete divergence-free condition, $(\operatorname{div}\boldsymbol{u}_{h},q_{h})=0$ for all $q_{h}\in P_{h}$, does not yield a strong divergence-free velocity field (i.e., $\operatorname{div}\boldsymbol{u}_{h}=0$ pointwise) because $P_{h}$ is only a proper subspace of $L^{2}_{0}(\Omega)$. Schemes that achieve this pointwise property are called strong divergence-free. In such schemes, one can choose an appropriate interpolant $\boldsymbol{u}_{I}$ such that $\operatorname{div}(\boldsymbol{u}_{I}-\boldsymbol{u}_{h})=0$ to decouple the velocity error from the pressure error via Galerkin orthogonality:
(3) $(\nabla(\boldsymbol{u}-\boldsymbol{u}_{h}),\nabla(\boldsymbol{u}_{I}-\boldsymbol{u}_{h}))=(p-p_{h},\operatorname{div}(\boldsymbol{u}_{I}-\boldsymbol{u}_{h}))=0.$
If the scheme is not strong divergence-free, the term $(p,\operatorname{div}(\boldsymbol{u}_{I}-\boldsymbol{u}_{h}))$ does not generally vanish, introducing a pressure dependency into the velocity error. When it does vanish, we obtain the bound $\|\nabla(\boldsymbol{u}-\boldsymbol{u}_{h})\|\leq\|\nabla(\boldsymbol{u}-\boldsymbol{u}_{I})\|$, which depends only on the regularity of the velocity. This decoupling yields a pressure-robust error estimate.
@@
Several methods provide strong divergence-free velocity approximations. These include the Scott–Vogelius element [50] (which requires specific mesh conditions), smooth finite elements [57, 18, 19], conforming pairs on split meshes [7, 56, 23, 29, 33, 39], and conforming rational functions [34, 35]. Low-order nonconforming pairs [45, 51, 55] are also used.
A natural way to enforce the strong divergence-free condition is to employ $H(\operatorname{div})$-conforming velocity spaces together with discontinuous polynomial pressure spaces. For example, using the Raviart–Thomas ($\mathrm{RT}_{k}$) [49, 47] or Brezzi–Douglas–Marini ($\mathrm{BDM}_{k}$) [13, 48] spaces for velocity (with $\boldsymbol{u}\cdot\boldsymbol{n}=0$ on $\partial\Omega$) paired with the discontinuous space $\mathbb{P}_{\ell}(\mathcal{T}_{h})$ for pressure ensures LBB stability and yields a strongly divergence-free velocity field.
However, because the discrete velocity space is not a subspace of $H^{1}(\Omega;\mathbb{R}^{d})$, the vector Laplacian $\Delta\boldsymbol{u}$ lacks a standard $H^{1}$ weak formulation.
To address this, the vorticity–velocity–pressure formulation [27, 28] introduces the vorticity $\boldsymbol{\omega}=\operatorname{curl}\boldsymbol{u}$ and treats the Laplacian as the Hodge Laplacian: $\Delta\boldsymbol{u}=\operatorname{curl}\operatorname{curl}\boldsymbol{u}-\nabla\operatorname{div}\boldsymbol{u}$. Yet, when enforcing no-slip boundary conditions ($\boldsymbol{u}=\boldsymbol{0}$), this formulation suffers from reduced stability and suboptimal convergence rates due to the absence of a proper Hilbert complex [4]. Symmetrical meshes can partially recover these rates, as seen in triangular MAC (TMAC) schemes [22], which lump mass to eliminate vorticity.
Alternatively, the mass-conserving mixed stress (MCS) formulation [31, 32] embeds $H(\operatorname{div})$ velocities in a pseudostress–velocity–pressure system. The finite element space for the stress is tangential–normal continuous. While standard stress formulations yield $\mathcal{O}(h^{k})$ convergence [32], enriching the stress space allows the $\mathrm{RT}_{k}$$\mathrm{P}_{k}$ pair to achieve $\mathcal{O}(h^{k+1})$ convergence, with requirement $k\geq 1$ [31]. A sophisticated inf-sup condition is proved for tangential–normal continuous traceless tensors.
Inspired by the MCS framework, we use tangential–normal continuous elements, $\Sigma_{k}^{\operatorname{tn}}$, to discretize the vector Laplacian. Because $\Sigma_{k}^{\operatorname{tn}}$ is not $H(\operatorname{div})$-conforming, we define a weak divergence operator, $\operatorname{div}_{w}$, whose adjoint is the weak deviatoric gradient, $\operatorname{dev}\operatorname{grad}_{w}$, acting on the velocity element.
The viability of this method depends on establishing an inf-sup condition for $\operatorname{div}_{w}$ over the stress–velocity pair. This presents a structural tension: the velocity space must be large enough (e.g., $\mathrm{RT}_{k}$) to maintain the strong divergence-free property—which is necessary to decouple the stress and velocity approximation errors—but small enough (e.g., $\mathrm{BDM}_{k}$) to ensure that $\operatorname{div}_{w}\Sigma_{k}^{\operatorname{tn}}$ remains surjective. We resolve this conflict by establishing the inf-sup condition directly on the divergence-free subspace:
(4) $\inf_{\boldsymbol{v}_{h}\in\mathrm{BDM}_{k}\cap\ker(\operatorname{div})}\sup_{\boldsymbol{\tau}_{h}\in\Sigma_{k}^{\operatorname{tn}}}\frac{(\operatorname{div}_{w}\boldsymbol{\tau}_{h},\boldsymbol{v}_{h})_{0,h}}{\|\boldsymbol{\tau}_{h}\|_{\operatorname{div}_{w}}\,\|\boldsymbol{v}_{h}\|}=\alpha>0,$
leveraging the crucial algebraic identity $\mathrm{BDM}_{k}\cap\ker(\operatorname{div})=\mathrm{RT}_{k}\cap\ker(\operatorname{div})$. Space $\mathrm{RT}_{k}$ is an enrichment of $\mathrm{BDM}_{k}$ to enlarge the range of the divergence operator. Furthermore, we establish the strong divergence-free property: if $(\operatorname{div}_{w}\boldsymbol{\tau}_{h},\boldsymbol{v}_{h})_{0,h}=0$ for all $\boldsymbol{v}_{h}\in\mathrm{BDM}_{k}\cap\ker(\operatorname{div})$, then in the distributional sense:
$(\operatorname{div}\boldsymbol{\tau}_{h},\boldsymbol{v})=0\qquad\forall\,\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d})\cap\ker(\operatorname{div}).$
The constraint $\operatorname{div}\boldsymbol{u}_{h}=0$ is imposed by introducing a Lagrange multiplier, which has the physical interpretation of pressure. Specifically, let $\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$ denote the $\mathrm{RT}_{k}$ space when $\ell=k$, or the $\mathrm{BDM}_{k}$ space when $\ell=k-1$, satisfying the boundary condition $\boldsymbol{u}\cdot\boldsymbol{n}=0$ on $\partial\Omega$. The discrete problem seeks $(\boldsymbol{u}_{h},p_{h})\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathbb{P}_{\ell}(\mathcal{T}_{h})$ for $k\geq 0$ and $\ell\in\{k,k-1\}$ such that:
@@
(5) $\begin{cases}a_{h}(\boldsymbol{u}_{h},\boldsymbol{v}_{h})+b(\boldsymbol{v}_{h},p_{h})=(\boldsymbol{f},\boldsymbol{v}_{h})&\text{for all }\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}},\\ b(\boldsymbol{u}_{h},q_{h})=0&\text{for all }q_{h}\in\mathbb{P}_{\ell}(\mathcal{T}_{h}),\end{cases}$
where $a_{h}(\boldsymbol{u}_{h},\boldsymbol{v}_{h}):=(\operatorname{dev}\operatorname{grad}_{w}\boldsymbol{u}_{h},\operatorname{dev}\operatorname{grad}_{w}\boldsymbol{v}_{h})$ and $b(\boldsymbol{v}_{h},q_{h}):=(\operatorname{div}\boldsymbol{v}_{h},q_{h})$. Section 2 details the precise definitions of these spaces and the $\operatorname{dev}\operatorname{grad}_{w}$ operator, while Section 3 presents a hybridized version of (5). The boundary condition of the tangential component $\boldsymbol{u}\cdot\boldsymbol{t}=0$ on $\partial\Omega$ is imposed weakly and can be replaced by $\boldsymbol{t}^{\intercal}\boldsymbol{\sigma}\boldsymbol{n}=0$ on $\partial\Omega$.
We emphasize that this method requires no additional stabilization. The variational formulation is not augmented by mesh-dependent penalty parameters or grad-div stabilization terms. Instead, stability follows directly from the discrete inf-sup conditions for the $\operatorname{div}_{w}$ (4) and $\operatorname{div}$ operators (2). The weak deviatoric gradient, $\operatorname{dev}\operatorname{grad}_{w}$, naturally incorporates face contributions via elementwise integration by parts, yielding a consistent, parameter-free discretization that recovers the continuous identity for smooth functions.
Let $\boldsymbol{\sigma}=\operatorname{dev}\operatorname{grad}\boldsymbol{u}$ and $\boldsymbol{\sigma}_{h}=\operatorname{dev}\operatorname{grad}_{w}\boldsymbol{u}_{h}\in\Sigma_{k}^{\operatorname{tn}}$ denote the continuous and discrete stresses, respectively. Owing to the strong divergence-free property of $\operatorname{div}_{w}$, the stress error decouples from the velocity and pressure (using argument similar to (3)), achieving optimal $\mathcal{O}(h^{k+1})$ convergence in $L^{2}$. This yields the following estimate:
$\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|+\|\operatorname{dev}\operatorname{grad}_{w}(\boldsymbol{u}_{I}-\boldsymbol{u}_{h})\|+\|Q_{\ell}p-p_{h}\|\lesssim h^{k+1}|\boldsymbol{\sigma}|_{k+1},$
where $\boldsymbol{u}_{I}$ is an interpolant of $\boldsymbol{u}$ into $\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$, and $Q_{\ell}$ is the $L^{2}$-orthogonal projection onto $\mathbb{P}_{\ell}(\mathcal{T}_{h})$. Because $h^{k+1}$ exceeds the standard approximation capabilities of the chosen velocity and pressure spaces, both the velocity error $\|\operatorname{dev}\operatorname{grad}_{w}(\boldsymbol{u}_{I}-\boldsymbol{u}_{h})\|$ and the pressure error $\|Q_{\ell}p-p_{h}\|$ are superconvergent. This structural superconvergence does not require mesh symmetry—unlike TMAC [22]—and it facilitates higher-order velocity post-processing.
Alternatively, discontinuous Galerkin (DG) methods can discretize the vector Laplacian [53, 43, 24]. However, the DG method in [53] requires a penalty term for stability, and the hybridizable DG (HDG) method in [24] uses $k$th-order polynomials for all variables, whereas for $\boldsymbol{u}_{h}\in\mathrm{BDM}_{k}$, it suffices to choose $p_{h}\in\mathbb{P}_{k-1}(\mathcal{T}_{h})$. Similarly, the divergence-free weak virtual element method (VEM) [21, 9, 54, 41] generally requires stabilization for the projected discrete gradient and relies on the full $\mathrm{BDM}_{k}$ space; using the incomplete $\mathrm{RT}_{k}$ space degrades convergence. Consequently, their energy-norm velocity estimates are optimal, but not superconvergent.
Our framework extends to the lowest-order case $\ell=0$, which the MCS framework does not address [31, 32]. Specifically, the $(k,\ell)=(0,0)$ configuration (illustrated in Fig. 1) generalizes the classical MAC scheme [36] to unstructured triangulations. It achieves the first-order convergence without the accuracy loss typical of vorticity-based formulations. Similarly, the $\mathrm{BDM}_{1}$$\mathbb{P}_{0}$ pair ($(k,\ell)=(1,0)$) provides a second-order scheme.
Figure 1. Lowest-order configuration $(k,\ell)=(0,0)$. The velocity–pressure pair is ${\rm RT}_{0}$$\mathbb{P}_{0}$, and the stress is a piecewise constant traceless matrix with tangential–normal continuity.
@@
We can further hybridize (5) by relaxing the normal continuity of the velocity, which permits the choice $(k,\ell)=(0,-1)$. Here, the velocity is approximated by piecewise constant vectors, and normal continuity is enforced via a nonconforming linear pressure space. This configuration maintains the first-order convergence.
[…]
Although $H(\operatorname{div})$-conforming velocity fields naturally enforce mass conservation, the primary challenge remains the consistent discretization of the vector Laplacian. We address this by developing a distributional framework in Section 2 and establishing the inf-sup condition for the stress–velocity pairs. In Section 3, we formulate the mixed finite element methods for the Stokes equations by introducing the pressure as a Lagrange multiplier to enforce the divergence-free constraint. We prove the stability of $\operatorname{div}_{w}$ onto $\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$, including the divergence-free subspace $\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cap\ker(\operatorname{div})$, and establish optimal-order error estimates. Furthermore, Section 4 explores equivalent discrete formulations, including virtual element and pseudostress methods. Finally, Section 5 provides numerical experiments to verify our theoretical findings, and Section 6 offers concluding remarks.